# Reasoning Under Uncertainty: Exploring Probabilistic Reasoning Capabilities of LLMs

### v1

@@ Appendix D Independence of Generated Samples
To investigate this phenomenon more systematically, we conducted additional analyses using two statistical methods: (1) the Durbin–Watson (DW) statistic [Durbin and Watson, 1951] to detect autocorrelation, and (2) transition matrix entropy [Shannon, 1948] to evaluate diversity in transition patterns. Samples are classified as non-independent based on the aggregated results of these two measures. Table 3 reports the corresponding statistics for samples generated by each model.
Appendix E Evaluation on Real-world Data
@@
To validate our findings on real-world data, we conducted experiments on the Mushroom dataset [mus, 1981], focusing on three categorical features and a subset of 80 samples. The resulting joint distribution is highly skewed, with a large proportion of zero-frequency outcomes, making the task particularly challenging. Table 4 summarizes the observed outcomes and their frequencies in this subset. The experimental results, detailed in Table 5, show that larger models, as well as distilled variants, consistently outperform smaller models across all task configurations.
Appendix F Effect of Sample Count in Context
@@
Appendix G Access to Code Interpreter

### v2

@@ Appendix E Independence of Generated Samples
To investigate this phenomenon more systematically, we conducted additional analyses using two statistical methods: (1) the Durbin–Watson (DW) statistic [8] to detect autocorrelation, and (2) transition matrix entropy [32] to evaluate diversity in transition patterns. Samples are classified as non-independent based on the aggregated results of these two measures. Table 4 reports the corresponding statistics for samples generated by each model.
Appendix F Evaluation on Real-world Data
To validate our findings on real-world data, we conducted experiments on the Mushroom dataset [1], focusing on three categorical features and a subset of 80 samples. The resulting joint distribution is highly skewed, with a large proportion of zero-frequency outcomes, making the task particularly challenging. Table 5 summarizes the observed outcomes and their frequencies in this subset. The experimental results, detailed in Table 6, show that larger models, as well as distilled variants, consistently outperform smaller models across all task configurations.
@@ Appendix F Evaluation on Real-world Data
Appendix G Effect of Sample Count in Context
@@
Appendix H Access to Code Interpreter